OpenStudy (anonymous):

You originally draw a design for an art contest on a 3 in. x 5 in. card. The second phase of the contest requires the drawing to be transferred to an 8.5 in x 11 in. standard sheet of paper and utilize as much of the space on the paper as possible. You determine that the largest size one of the dimensions of your drawing can be is 10.5 in. What is the length of the other dimension if the two drawings are similar? Type your exact answer in the blank without the units, and round to the nearest tenths. PLEASE HELP

OpenStudy (anonymous):

please hhelp ill medal

OpenStudy (jdoe0001):

you made a 3x5 drawing then you're asked to transfter to a standard paper 8 1/2 x 11 says the largest of the dimensions comes up to 10.5 well, the largest of a 3x5 is 5 :) so one can say \(\begin{array}{llll} &width&length \\\hline\\ 3\times 5&3&5\\ 8.5\times 11\qquad &x&10.5 \end{array}\implies \cfrac{3}{x}=\cfrac{5}{10.5}\impliedby \textit{solve for "x"}\)